Suppose the call money rate is 6.8 percent, and you pay a spread of 1.9 percent over that. You buy 600 shares at $78 per share with an initial margin of 60 percent. One year later, the stock is selling for $84 per share and you close out your position. What is your return assuming no dividends are paid? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Initial equity = (600 × $78)(0.60) = $28,080

Amount borrowed = (600 × $78)(1 – 0.60) = $18,720

Interest = $18,720 × (0.0680 + 0.0190) = $18,720 × 0.0870 = $1,628.64

Proceeds from sale = 600 × $84 = $50,400

Dollar return = $50,400 - $18,720 - $1,628.64 - $28,080 = $1,971.36

Rate of return = $1,971.36 / $28,080 = 0.0702 or 7.02%